An auto mobile tire is inflated to be 0398 as decimal plus 32 psig pressure at 50 degree Fahrenheit after being driven the temperature rises to 75 degree Fahrenheit. Assuming that the volume remains constant, the final gauge pressure is?


1
Expert's answer
2022-02-15T16:17:07-0500

We know that

Idial gas equation

"T_1=50°F=283K\\T_2=75°F=297F"

"PV=nRT"

"\frac{P}{T}=\frac{R}{V}"

"\frac{P_1}{T_1}=\frac{P_2}{T_2}"

"P_2=\frac{P_1T_2}{T_1}"

"P_2=\frac{32\times297}{283}=33.58psia"